# Prompt Optimization with EASE? Efficient Ordering-aware Automated Selection of Exemplars

### v1

@@ 5.2 Empirical validation of the hypothesis through progressive finetuning
We hypothesize that the effect of exemplar selection diminishes as the model gains more knowledge about the task. To gain insights on this hypothesis, we study an open-source white-source Vicuna model instead of the black-box GPT-3.5 model that is only accessible through API. We progressively finetune the while-box language model, and examine whether the extent of finetuning on a specific task diminishes the importance of in-context exemplars when prompted.
@@ 5.3 New families of “out-of-distribution” tasks that emphasize in-context reasoning
Rule-based tasks. Given our insight on the impact of the model’s existing knowledge about the task (Sec. 5.2), we propose rule-based tasks that contain novel rules that the LLM has not learned before. A key characteristic of these tasks is that the model has to extract the underlying relationships among the provided in-context exemplars and directly use the relationship for test-time inferences. For example, we construct the linear regression (LR) task where the input takes the form demonstrated in Example B.1(see App. B.1). The underlying relationship in this example is $y=ax+b$, with $a=-4$ and $b=6$ in this specific Example B.1. Without further instructions, the model is supposed to rely on the provided in-context exemplars to implicitly infer the regression task, recover the coefficients $a,b$ for linear regression, and then directly apply it to the test sample (e.g., for test input $117$, compute $-4\times 117+6=-462$ and output $462$). The results are in Tab. 2. For the clean dataset (i.e., 0% noise), EASE outperforms the most competitive baseline by 10% in absolute accuracy. Note that EASE has greater advantages in settings with noisy data, which will be discussed later in this section.
@@
Remapped label tasks. By remapping the labels of existing classification tasks to new ones, we construct novel tasks that are against the model’s existing knowledge. In the commonly used AG News dataset, the news articles are classified into four categories: “World”, “Sports”, “Business” and “Sci/Tech”. We construct a remapped dataset AG News Remap such that “World” news is now labeled as “Sports” news, “Sports” is now labeled as “Business”, etc. This is against the LLM’s knowledge since the output now does not correspond to the context descriptions of the input (i.e., news articles). So, the LLM has to learn these remapping rules from the in-context exemplars. Similarly, we construct another dataset SST5 Reversed by reversing the labels of the sentiment analysis dataset SST5, such that “very negative” labels are swapped with “very positive” labels, and “negative” labels are swapped with “positive” labels. The results for these novel remapped label tasks are also in Tab. 2, which shows superior performances of EASE over baselines by about 10% in absolute accuracy.
@@
Note that the conclusions on test accuracy results (in App. C.10) are consistent with the main text. These tasks above represent new families of tasks that contain novel knowledge for LLMs. We show through comprehensive experiments that exemplar selection is important and useful in practice, especially for novel downstream task adaptations. Also, the noisy datasets considered here align with real-world scenarios, which typically contain noises from observation, labeling error, corruption, etc.
@@
Rule-based tasks LR 0% 31.7$\pm$1.4 38.3$\pm$2.7 50.0$\pm$0.0 71.7$\pm$1.4 70.0$\pm$0.0 36.7$\pm$1.4 56.7$\pm$5.9 61.7$\pm$1.4 66.7$\pm$1.4 81.7$\pm$3.6
10% 8.3$\pm$1.4 36.7$\pm$1.4 48.3$\pm$1.4 61.7$\pm$1.4 61.7$\pm$1.4 0.0$\pm$0.0 58.3$\pm$3.6 60.0$\pm$0.0 65.0$\pm$2.4 73.3$\pm$3.6
30% 10.0$\pm$0.0 28.3$\pm$1.4 46.7$\pm$1.4 63.3$\pm$1.4 60.0$\pm$0.0 40.0$\pm$2.4 35.0$\pm$2.4 53.3$\pm$1.4 50.0$\pm$0.0 78.3$\pm$1.4
50% 0.0$\pm$0.0 38.3$\pm$1.4 45.0$\pm$0.0 65.0$\pm$0.0 53.3$\pm$1.4 0.0$\pm$0.0 53.3$\pm$1.4 46.7$\pm$1.4 45.0$\pm$0.0 71.7$\pm$2.7
70% 0.0$\pm$0.0 55.0$\pm$0.0 38.3$\pm$2.7 65.0$\pm$0.0 50.0$\pm$0.0 26.7$\pm$5.4 30.0$\pm$4.7 33.3$\pm$1.4 33.3$\pm$1.4 66.7$\pm$3.6
90% 0.0$\pm$0.0 21.7$\pm$1.4 26.7$\pm$1.4 46.7$\pm$1.4 3.3$\pm$1.4 0.0$\pm$0.0 6.7$\pm$2.7 8.3$\pm$1.4 15.0$\pm$0.0 53.3$\pm$2.7
LP-variant 0% 48.3$\pm$2.7 40.0$\pm$2.4 41.7$\pm$1.4 65.0$\pm$0.0 58.3$\pm$1.4 30.0$\pm$0.0 61.7$\pm$1.4 75.0$\pm$2.4 71.7$\pm$1.4 75.0$\pm$0.0
10% 0.0$\pm$0.0 36.7$\pm$1.4 40.0$\pm$0.0 63.3$\pm$2.7 60.0$\pm$0.0 36.7$\pm$2.7 65.0$\pm$2.4 70.0$\pm$2.4 73.3$\pm$1.4 75.0$\pm$2.4
30% 0.0$\pm$0.0 48.3$\pm$2.7 40.0$\pm$2.4 60.0$\pm$0.0 55.0$\pm$0.0 40.0$\pm$7.1 53.3$\pm$4.9 65.0$\pm$2.4 65.0$\pm$0.0 73.3$\pm$1.4
50% 0.0$\pm$0.0 65.0$\pm$0.0 35.0$\pm$2.4 63.3$\pm$2.7 60.0$\pm$0.0 38.3$\pm$3.6 48.3$\pm$3.6 61.7$\pm$1.4 65.0$\pm$0.0 76.7$\pm$2.7
70% 0.0$\pm$0.0 46.7$\pm$2.7 35.0$\pm$0.0 70.0$\pm$0.0 60.0$\pm$0.0 25.0$\pm$8.2 60.0$\pm$4.1 56.7$\pm$1.4 56.7$\pm$1.4 75.0$\pm$0.0
90% 0.0$\pm$0.0 35.0$\pm$2.4 50.0$\pm$0.0 65.0$\pm$2.4 0.0$\pm$0.0 30.0$\pm$12.5 50.0$\pm$2.4 38.3$\pm$1.4 55.0$\pm$2.4 63.3$\pm$1.4
Re-mapped label tasks AG News Remap 0% 20.0$\pm$2.4 15.0$\pm$0.0 26.7$\pm$1.4 43.3$\pm$1.4 43.3$\pm$2.7 5.0$\pm$2.4 25.0$\pm$4.1 40.0$\pm$0.0 40.0$\pm$0.0 53.3$\pm$3.6
10% 5.0$\pm$0.0 15.0$\pm$0.0 15.0$\pm$0.0 41.7$\pm$1.4 38.3$\pm$1.4 3.3$\pm$1.4 26.7$\pm$2.7 36.7$\pm$1.4 40.0$\pm$0.0 56.7$\pm$2.7
30% 10.0$\pm$0.0 5.0$\pm$0.0 5.0$\pm$0.0 40.0$\pm$0.0 36.7$\pm$1.4 1.7$\pm$1.4 10.0$\pm$0.0 40.0$\pm$0.0 43.3$\pm$1.4 51.7$\pm$1.4
50% 5.0$\pm$0.0 10.0$\pm$0.0 5.0$\pm$0.0 43.3$\pm$1.4 35.0$\pm$0.0 3.3$\pm$1.4 20.0$\pm$4.1 35.0$\pm$0.0 35.0$\pm$0.0 56.7$\pm$1.4
70% 5.0$\pm$0.0 25.0$\pm$0.0 8.3$\pm$1.4 50.0$\pm$0.0 35.0$\pm$0.0 1.7$\pm$1.4 11.7$\pm$5.4 38.3$\pm$1.4 46.7$\pm$1.4 51.7$\pm$1.4
90% 5.0$\pm$0.0 18.3$\pm$1.4 5.0$\pm$0.0 40.0$\pm$0.0 10.0$\pm$0.0 15.0$\pm$6.2 35.0$\pm$0.0 35.0$\pm$0.0 41.7$\pm$1.4 55.0$\pm$2.4
SST5 Reverse 0% 20.0$\pm$0.0 10.0$\pm$0.0 13.3$\pm$1.4 40.0$\pm$0.0 40.0$\pm$0.0 15.0$\pm$2.4 33.3$\pm$5.4 35.0$\pm$2.4 40.0$\pm$0.0 50.0$\pm$0.0
10% 16.7$\pm$1.4 10.0$\pm$0.0 15.0$\pm$0.0 48.3$\pm$1.4 40.0$\pm$0.0 13.3$\pm$2.7 23.3$\pm$5.4 33.3$\pm$2.7 40.0$\pm$0.0 50.0$\pm$0.0
30% 23.3$\pm$1.4 6.7$\pm$1.4 25.0$\pm$2.4 40.0$\pm$0.0 40.0$\pm$0.0 21.7$\pm$3.6 26.7$\pm$1.4 30.0$\pm$0.0 31.7$\pm$1.4 41.7$\pm$3.6
50% 21.7$\pm$1.4 15.0$\pm$0.0 15.0$\pm$0.0 43.3$\pm$1.4 33.3$\pm$1.4 21.7$\pm$1.4 23.3$\pm$1.4 28.3$\pm$1.4 30.0$\pm$0.0 43.3$\pm$1.4
70% 25.0$\pm$0.0 23.3$\pm$1.4 23.3$\pm$1.4 40.0$\pm$0.0 30.0$\pm$0.0 20.0$\pm$2.4 25.0$\pm$2.4 36.7$\pm$1.4 36.7$\pm$1.4 45.0$\pm$2.4
90% 20.0$\pm$0.0 15.0$\pm$2.4 20.0$\pm$0.0 30.0$\pm$0.0 30.0$\pm$0.0 13.3$\pm$2.7 21.7$\pm$1.4 30.0$\pm$0.0 30.0$\pm$0.0 31.7$\pm$1.4
@@ 5.4 Jointly optimizing instruction and exemplars
To the best of our knowledge, no prior work can jointly optimize instruction and exemplars. For a fair comparison, we do not include an instruction when comparing with other exemplar selection baselines in earlier sections. As EASE is readily extended to find the optimal combination of instruction and exemplars (Sec. 4.3), we show the benefits of joint optimization in this section. According to Tab. 3, jointly optimizing these two essential components of a prompt significantly improves the performance for some difficult tasks (marked with red arrows $\uparrow$). This improvement is attributed to the ability of the optimized instruction to significantly reinforce the information captured in the corresponding exemplars. For example, an automatically optimized instruction “identify and list the animals from the given words” complements the exemplars in the taxonomy animal task. Therefore, our joint optimization of instruction and exemplars presents a fully automated pipeline for prompt optimization and achieves impressive practical performances.
@@ 5.5 Ablation studies
Combination with retrieval-based methods to handle larger sets of exemplars. Applying EASE to scenarios where the size $n$ of the set of exemplars $D$ is very large may lead to performance degradation. This is because the space $\Omega$ of exemplar sequences becomes excessively large, which cannot be sufficiently explored without significantly increasing the computation (i.e., with the size of $Q_{t}$ being fixed). To resolve this issue, a natural idea is to first filter the large pool of data exemplars to eliminate those that are less relevant to the task. To this end, we propose to first use retrieval-based methods to select the exemplars that are more relevant to the task, and then run our EASE using this refined smaller set of exemplars. Specifically, we use a cosine similarity retriever and perform exemplar selection on $D$ with a size $n$ as large as 100000. As shown in Tab. 4, when the size $n$ of the exemplar set is large, combining EASE with retrieval gives better performances than directly running EASE.
Table 3: Average accuracy $\pm$ s.e. for EASE with and without jointly optimized instructions. We removed tasks with 100% accuracy. The full results are in App C, Tab. 6.
@@
auto_categorization 30.0$\pm$0.0 46.7$\pm$4.9 16.7 $\uparrow$
@@
object_counting 73.3$\pm$1.4 75.0$\pm$0.0 1.7 $\uparrow$
orthography_starts_with 78.3$\pm$1.4 81.7$\pm$1.4 3.3 $\uparrow$
rhymes 100.0$\pm$0.0 91.7$\pm$3.6 -8.3 $\downarrow$
second_word_letter 50.0$\pm$0.0 100.0$\pm$0.0 50.0 $\uparrow$
sentence_similarity 56.7$\pm$1.4 56.7$\pm$1.4 0.0 $\circ$
synonyms 30.0$\pm$0.0 30.0$\pm$0.0 0.0 $\circ$
taxonomy_animal 88.3$\pm$2.7 100.0$\pm$0.0 11.7 $\uparrow$
@@
translation_en-es 100.0$\pm$0.0 100.0$\pm$0.0 0.0 $\circ$
translation_en-fr 88.3$\pm$1.4 85.0$\pm$0.0 -3.3 $\downarrow$
word_sorting 91.7$\pm$1.4 91.7$\pm$1.4 0.0 $\circ$
word_unscrambling 78.3$\pm$2.7 80.0$\pm$0.0 1.7 $\uparrow$
linear_4_10_noisy 73.3$\pm$3.6 41.7$\pm$9.5 -31.7 $\downarrow$
LP-variant (10% noise) 75.0$\pm$2.4 85.0$\pm$2.4 10.0 $\uparrow$
AG News Remap (10% noise) 56.7$\pm$2.7 65.0$\pm$0.0 8.3 $\uparrow$
SST5 Reverse (10% noise) 50.0$\pm$0.0 50.0$\pm$0.0 0.0 $\circ$
@@
Further ablations. We defer further ablation studies to App. C, including those exploring (a) speedups from OT, (b) a larger $k$, (c) benefits of using an ordering-aware embedding, (d) different black-box LLMs, (e) other embedding models, and (f) asking GPT to directly select exemplars.

### v2

@@ 4.2 Empirical validation of the hypothesis through progressive finetuning
We hypothesize that the effect of exemplar selection diminishes as the model gains more knowledge about the task. To gain insights on this hypothesis, we study an open-source white-box Vicuna model instead of the black-box GPT-3.5 model that is only accessible through API. We progressively finetune the while-box language model, and examine whether the extent of finetuning on a specific task diminishes the importance of in-context exemplars when prompted.
@@ 4.3 New families of “out-of-distribution” tasks that emphasize in-context reasoning
Rule-based tasks. Given our insight on the impact of the model’s existing knowledge about the task (Sec. 4.2), we propose rule-based tasks that contain novel rules that the LLM has not learned before. A key characteristic of these tasks is that the model has to extract the underlying relationships among the provided in-context exemplars and directly use the relationship for test-time inferences. For example, we construct the linear regression (LR) task where the input takes the form demonstrated in Example B.1 (see App. B.1). The underlying relationship in this example is $y=ax+b$, with $a=-4$ and $b=6$ in this specific Example B.1. Without further instructions, the model is supposed to rely on the provided in-context exemplars to implicitly infer the regression task, recover the coefficients $a,b$ for linear regression, and then directly apply it to the test sample (e.g., for test input $117$, compute $-4\times 117+6=-462$ and output $-462$). The results are in Tab. 2. For the clean dataset (i.e., 0% noise), EASE outperforms the most competitive baseline by 8.3% in absolute accuracy. Note that EASE has greater advantages in settings with noisy data, which will be discussed later in this section.
@@
Remapped label tasks. By remapping the labels of existing classification tasks to new ones, we construct novel tasks that are against the model’s existing knowledge. In the commonly used AG News dataset, the news articles are classified into four categories: “World”, “Sports”, “Business” and “Sci/Tech”. We construct a remapped dataset AG News Remap such that “World” news is now labeled as “Sports” news, “Sports” is now labeled as “Business”, etc. This is against the LLM’s knowledge since the output now does not correspond to the context descriptions of the input (i.e., news articles). So, the LLM has to learn these remapping rules from the in-context exemplars. Similarly, we construct another dataset SST5 Reversed by reversing the labels of the sentiment analysis dataset SST5, such that “very negative” labels are swapped with “very positive” labels, and “negative” labels are swapped with “positive” labels. The results for these novel remapped label tasks are also in Tab. 2, which shows superior performances of EASE over baselines by 6.7%-10% in absolute accuracy.
@@
Note that the conclusions on test accuracy results (in App. C.15) are consistent with the main text. These tasks above represent new families of tasks that contain novel knowledge for LLMs. We show through comprehensive experiments that exemplar selection is important and useful in practice, especially for novel downstream task adaptations. Also, the noisy datasets considered here align with real-world scenarios, which typically contain noises from observation, labeling error, corruption, etc.
@@
Rule-based tasks LR 0% 31.7$\pm$1.7 38.3$\pm$3.3 50.0$\pm$0.0 71.7$\pm$1.7 70.0$\pm$0.0 36.7$\pm$1.4 56.7$\pm$7.3 61.7$\pm$1.7 66.7$\pm$1.7 80.0$\pm$2.9
10% 8.3$\pm$1.7 36.7$\pm$1.7 48.3$\pm$1.7 61.7$\pm$1.7 61.7$\pm$1.7 0.0$\pm$0.0 58.3$\pm$4.4 60.0$\pm$0.0 65.0$\pm$2.9 73.3$\pm$1.7
30% 10.0$\pm$0.0 28.3$\pm$1.7 46.7$\pm$1.7 63.3$\pm$1.7 60.0$\pm$0.0 40.0$\pm$2.4 35.0$\pm$2.9 53.3$\pm$1.7 50.0$\pm$0.0 76.7$\pm$1.7
50% 0.0$\pm$0.0 38.3$\pm$1.7 45.0$\pm$0.0 65.0$\pm$0.0 53.3$\pm$1.7 0.0$\pm$0.0 53.3$\pm$1.7 46.7$\pm$1.7 45.0$\pm$0.0 78.3$\pm$4.4
70% 0.0$\pm$0.0 55.0$\pm$0.0 38.3$\pm$3.3 65.0$\pm$0.0 50.0$\pm$0.0 26.7$\pm$5.4 30.0$\pm$5.8 33.3$\pm$1.7 33.3$\pm$1.7 66.7$\pm$1.7
90% 0.0$\pm$0.0 21.7$\pm$1.7 26.7$\pm$1.7 46.7$\pm$1.7 3.3$\pm$1.7 0.0$\pm$0.0 6.7$\pm$3.3 8.3$\pm$1.7 15.0$\pm$0.0 53.3$\pm$1.7
LP-variant 0% 48.3$\pm$3.3 40.0$\pm$2.9 41.7$\pm$1.7 65.0$\pm$0.0 58.3$\pm$1.7 30.0$\pm$0.0 61.7$\pm$1.7 75.0$\pm$2.9 71.7$\pm$1.7 75.0$\pm$0.0
10% 0.0$\pm$0.0 36.7$\pm$1.7 40.0$\pm$0.0 63.3$\pm$3.3 60.0$\pm$0.0 36.7$\pm$2.7 65.0$\pm$2.9 70.0$\pm$2.9 73.3$\pm$1.7 80.0$\pm$2.9
30% 0.0$\pm$0.0 48.3$\pm$3.3 40.0$\pm$2.9 60.0$\pm$0.0 55.0$\pm$0.0 40.0$\pm$7.1 53.3$\pm$6.0 65.0$\pm$2.9 65.0$\pm$0.0 75.0$\pm$0.0
50% 0.0$\pm$0.0 65.0$\pm$0.0 35.0$\pm$2.9 63.3$\pm$3.3 60.0$\pm$0.0 38.3$\pm$3.6 48.3$\pm$4.4 61.7$\pm$1.7 65.0$\pm$0.0 78.3$\pm$1.7
70% 0.0$\pm$0.0 46.7$\pm$3.3 35.0$\pm$0.0 70.0$\pm$0.0 60.0$\pm$0.0 25.0$\pm$8.2 60.0$\pm$5.0 56.7$\pm$1.7 56.7$\pm$1.7 71.7$\pm$1.7
90% 0.0$\pm$0.0 35.0$\pm$2.9 50.0$\pm$0.0 65.0$\pm$2.9 0.0$\pm$0.0 30.0$\pm$12.5 50.0$\pm$2.9 38.3$\pm$1.7 55.0$\pm$2.9 66.7$\pm$1.7
Re-mapped label tasks AG News Remap 0% 20.0$\pm$2.9 15.0$\pm$0.0 26.7$\pm$1.7 43.3$\pm$1.7 43.3$\pm$3.3 5.0$\pm$2.4 25.0$\pm$5.0 40.0$\pm$0.0 40.0$\pm$0.0 50.0$\pm$0.0
10% 5.0$\pm$0.0 15.0$\pm$0.0 15.0$\pm$0.0 41.7$\pm$1.7 38.3$\pm$1.7 3.3$\pm$1.4 26.7$\pm$3.3 36.7$\pm$1.7 40.0$\pm$0.0 51.7$\pm$1.7
30% 10.0$\pm$0.0 5.0$\pm$0.0 5.0$\pm$0.0 40.0$\pm$0.0 36.7$\pm$1.7 1.7$\pm$1.4 10.0$\pm$0.0 40.0$\pm$0.0 43.3$\pm$1.7 55.0$\pm$0.0
50% 5.0$\pm$0.0 10.0$\pm$0.0 5.0$\pm$0.0 43.3$\pm$1.7 35.0$\pm$0.0 3.3$\pm$1.4 20.0$\pm$5.0 35.0$\pm$0.0 35.0$\pm$0.0 55.0$\pm$2.9
70% 5.0$\pm$0.0 25.0$\pm$0.0 8.3$\pm$1.7 50.0$\pm$0.0 35.0$\pm$0.0 1.7$\pm$1.4 11.7$\pm$6.7 38.3$\pm$1.7 46.7$\pm$1.7 58.3$\pm$3.3
90% 5.0$\pm$0.0 18.3$\pm$1.7 5.0$\pm$0.0 40.0$\pm$0.0 10.0$\pm$0.0 15.0$\pm$6.2 35.0$\pm$0.0 35.0$\pm$0.0 41.7$\pm$1.7 53.3$\pm$1.7
SST5 Reverse 0% 20.0$\pm$0.0 10.0$\pm$0.0 13.3$\pm$1.7 40.0$\pm$0.0 40.0$\pm$0.0 15.0$\pm$2.4 33.3$\pm$6.7 35.0$\pm$2.9 40.0$\pm$0.0 50.0$\pm$2.9
10% 16.7$\pm$1.7 10.0$\pm$0.0 15.0$\pm$0.0 48.3$\pm$1.7 40.0$\pm$0.0 13.3$\pm$2.7 23.3$\pm$6.7 33.3$\pm$3.3 40.0$\pm$0.0 48.3$\pm$1.7
30% 23.3$\pm$1.7 6.7$\pm$1.7 25.0$\pm$2.9 40.0$\pm$0.0 40.0$\pm$0.0 21.7$\pm$3.6 26.7$\pm$1.7 30.0$\pm$0.0 31.7$\pm$1.7 46.7$\pm$3.3
50% 21.7$\pm$1.7 15.0$\pm$0.0 15.0$\pm$0.0 43.3$\pm$1.7 33.3$\pm$1.7 21.7$\pm$1.4 23.3$\pm$1.7 28.3$\pm$1.7 30.0$\pm$0.0 46.7$\pm$3.3
70% 25.0$\pm$0.0 23.3$\pm$1.7 23.3$\pm$1.7 40.0$\pm$0.0 30.0$\pm$0.0 20.0$\pm$2.4 25.0$\pm$2.9 36.7$\pm$1.7 36.7$\pm$1.7 45.0$\pm$5.0
90% 20.0$\pm$0.0 15.0$\pm$2.9 20.0$\pm$0.0 30.0$\pm$0.0 30.0$\pm$0.0 13.3$\pm$2.7 21.7$\pm$1.7 30.0$\pm$0.0 30.0$\pm$0.0 31.7$\pm$1.7
@@ 4.4 Jointly optimizing instruction and exemplars
To the best of our knowledge, no prior work can jointly optimize instruction and exemplars. For a fair comparison, we do not include an instruction when comparing with other exemplar selection baselines in earlier sections. As EASE is readily extended to find the optimal combination of instruction and exemplars (Sec. 3.3), we show the benefits of joint optimization in this section. According to Tab. 3, jointly optimizing these two essential components of a prompt significantly improves the performance for most tasks (marked with red arrows $\uparrow$). This improvement is attributed to the ability of the optimized instruction to significantly reinforce the information captured in the corresponding exemplars. For example, an automatically optimized instruction “identify and list the animals from the given words” complements the exemplars in the taxonomy animal task. Therefore, our joint optimization of instruction and exemplars presents a fully automated pipeline for prompt optimization and achieves impressive practical performances.
@@ 4.5 Ablation studies
Combination with retrieval-based methods to handle larger sets of exemplars. Applying EASE to scenarios where the size $n$ of the set of exemplars $D$ is very large may lead to performance degradation. This is because the space $\Omega$ of exemplar sequences becomes excessively large, which cannot be sufficiently explored without significantly increasing the computation (i.e., with the size of $Q_{t}$ being fixed). To resolve this issue, a natural idea is to first filter the large pool of data exemplars to eliminate those that are less relevant to the task. This independent step to pre-filter candidates serves as a promising extension of our EASE to handle a large set of exemplars. To this end, we propose to first use retrieval-based methods to select the exemplars that are more relevant to the task, and then run our EASE using this refined smaller set of exemplars. Specifically, we use a cosine similarity retriever on exemplar embedding and perform exemplar selection on $D$ with a size $n$ as large as 100000. The query involves validation data exemplars $D_{V}$ and we retrieve from the corpus of training set exemplars $D$. For each validation data exemplar in $D_{V}$, we retrieve the most similar/relevant training set exemplars from $D$. Then, we combine them to form a smaller set of exemplars than $D$ and proceed with EASE using this reduced subset for efficiency. As shown in Tab. 4, when the size $n$ of the exemplar set is large, combining EASE with retrieval gives better performances than directly running EASE.
Table 3: Average accuracy $\pm$ s.e. for EASE with and without jointly optimized instructions. We removed tasks with 100% accuracy. The full results are in App. C, Tab. 6.
@@
auto_categorization 30.0$\pm$0.0 56.7$\pm$1.7 26.7 $\uparrow$
@@
object_counting 73.3$\pm$1.7 75.0$\pm$0.0 1.7 $\uparrow$
orthography_starts_with 80.0$\pm$0.0 81.7$\pm$1.7 1.7 $\uparrow$
second_word_letter 53.3$\pm$1.7 100.0$\pm$0.0 46.7 $\uparrow$
sentence_similarity 56.7$\pm$1.7 58.3$\pm$1.7 1.7 $\uparrow$
synonyms 30.0$\pm$0.0 31.7$\pm$1.7 1.7 $\uparrow$
taxonomy_animal 88.3$\pm$1.7 100.0$\pm$0.0 11.7 $\uparrow$
@@
translation_en-fr 88.3$\pm$1.7 85.0$\pm$0.0 -3.3 $\downarrow$
word_sorting 90.0$\pm$0.0 93.3$\pm$1.7 3.3 $\uparrow$
word_unscrambling 78.3$\pm$1.7 80.0$\pm$0.0 1.7 $\uparrow$
LR (10% noise) 73.3$\pm$1.7 45.0$\pm$15.0 -28.3 $\downarrow$
LP-variant (10% noise) 80.0$\pm$2.9 86.7$\pm$1.7 6.7 $\uparrow$
AG News Remap (10% noise) 51.7$\pm$1.7 65.0$\pm$0.0 13.3 $\uparrow$
SST5 Reverse (10% noise) 48.3$\pm$1.7 53.3$\pm$1.7 5.0 $\uparrow$
@@
Further ablations. We defer further ablation studies to App. C, including those exploring (a) speedups from OT, (b) a larger $k$ or a range of $k$, (c) a larger query budget $T$, (d) benefits of using an ordering-aware embedding, (e) different black-box LLMs, (f) other embedding models, (g) degree of exploration $\nu$, (h) asking GPT to directly select exemplars, and (i) additional experiments on more datasets including reasoning chains.
5 Related work
Instruction optimization. Another related direction for enhancing the performance of LLMs is instruction optimization. Focusing on instructions within the prompt, evolutionary algorithms and zeroth-order optimization algorithms are gaining popularity in refining the prompts for black-box LLMs [5, 10, 13, 16, 21, 47]. Specifically, some studies [10, 16, 21, 47] maintained a constant set of exemplars throughout the process of prompt optimization, whereas others [5, 9, 13, 34] ignored the consideration of exemplars altogether by adopting a zero-shot setting. It is therefore imperative to develop an effective exemplar selection method for black-box LLMs.